If 4g of hydrogen gas occupies 20L, what will be the volume of 33g of CO2 gas at the same conditions?
1
Expert's answer
2020-10-16T08:59:46-0400


We know that dihydrogen (H2) has 2 hydrogen atoms per molecule, so there will be 2 * 1g hydrogen per mole = 2g dihydrogen per mole. If there are 2g dyhydrogen per mole, and we have 4g, then we know 4g / 2g per mole = 2 mole dihydrogen in 4g. So n1=2mole

So asusal 33g Co2=33/44 mole Co2

=0.75mole

So,n2=0.75mole

Volume of hydrogen gas(V1)=20L

Volume of Co2(V2)=?


So from ideal gas equestion we can write

P1V1/n1T1=P2V2/n2T2

Here,the temperature and pressure are constant.so we can write this equation

V1/n1=V2/n2

20L/2mole=V2/0.75mole

V2=0.75mole*20L/2

V2=7.5L

Hence the volume of CO2 is 7.5L